starting from the same point, Biden starts walking eastward at 52 m/s while Trump starts running towards the south at 86 m/s. How fast is the distance between Biden and Trump increasing after 2 seconds? (Input your final answer on the box below without units of measure


1
Expert's answer
2021-04-14T14:13:46-0400

In 2secs Biden would have covered 104m


While Trump would have covered 172m


Using Pythagoras theorem

"x^2+y^2=z^2"

x=104 y=172 while z is the distance between them


"z=\sqrt{104^2+172^2}=201m"

differentiating the Pythagoras theorem we have "2x{dx \over dt} + 2y{dy \over dt}=2z{dz \over dt}"


"{dx \over dt}= 52m/s" "{dy \over dt}=86m/s" "{dx \over dt}= ?"


"2(104)(52) +2(172)(86)=2(201){dx \over dt}"


"{dz \over dt}=100.5m/s"